# Flux of Extragalactic Dark Matter in Direct Detection Experiments

### v1

@@ Abstract
We calculate the contribution of extragalactic dark matter to the local dark matter density and flux in the Milky Way. By analyzing the Galactic escape velocity as a function of direction, we establish the criterion for separating dark matter particles bound to the Milky Way and those originating from the Local Group environment. Our analysis reveals that approximately 25% of dark matter particles in the Solar neighborhood have an extragalactic origin, contributing nearly 38% of the total mass flux. The directional dependence of this extragalactic component shows significant anisotropy across the sky, with implications for direct detection experiments. We provide quantitative predictions for detection rates and signatures that could help identify the extragalactic dark matter component in current and future experiments.
@@ I Introduction
Dark matter constitutes about 85% of the matter content in our universe, yet its nature remains elusive. Past predictions for the detection rate of laboratory experiments focused on particles that are gravitationally-bound to the Milky Way  [1, 2, 3, 4].Here, we show that a significant component originates from beyond our Galaxy. Inclusion of the extragalactic contribution could impact the quantitative interpretation of direct detection experiments.
To identify extragalactic dark matter particles, we calculate the escape velocity as a function of direction across the sky. Particles exceeding this threshold cannot be gravitationally bound to the Milky Way and must have an extragalactic origin. This condition provides a method for distinguishing galactic from extragalactic particles; those moving faster than the local escape velocity in a given direction must originate beyond the Milky Way’s gravitational influence.
The extragalactic component primarily originates from the Local Group environment surrounding our Galaxy. The Local Group, dominated by the Milky Way, Andromeda and the intergalactic matter around them, contains of order $\sim 5\times 10^{12}M_{\odot}$ of mass, with the baryonic component representing approximately 17% of the total [5]. This intergalactic medium supplies high-velocity dark matter particles, potentially detectable in Earth-based experiments. Recent N-body simulations [6] have demonstrated that the highest velocity dark matter particles (v $>$ 600 km/s) in Milky Way-like halos predominantly originate from fast-moving substructure, particularly Large Magellanic Cloud analogs, rather than from the broader extragalactic environment. This finding suggests a multi-component environment of high-velocity dark matter sources that must be considered in detection experiments.
@@ II Galactic Escape Velocity
Figure 1: Minimum escape velocity as a function of direction in a spherical projection of Galactic coordinates within our model. The map displays significant angular variation, with the highest escape velocities ($>700$ km/s, yellow region) concentrated around galactic coordinates $(90^{\circ},0^{\circ})$ in the eastern hemisphere, while the lowest escape velocities ($<400$ km/s, dark purple) appear in the opposite direction. This anisotropy reflects the LSR motion and the asymmetric gravitational potential of the Milky Way.
@@ IV.1 Extragalactic Flux Analysis
Figure 4: The extragalactic dark matter flux in GeV/cm2/s (×$\times 10^{6}$), shows significant angular variation with the highest flux values ($\sim 2\times 10^{7}$ GeV/cm2/s, yellow regions) concentrated in the western galactic hemisphere and a pronounced minimum in the opposite direcction.
@@ IV.2 Implications for Detection
Our analysis indicates that the extragalactic component contributes $38\%$ of the total mass flux while representing only $25\%$ of the particle number density. This kinematic distinction suggests that energy-selective detection methods could preferentially sample the extragalactic population, even without explicit directional sensitivity. Specifically, focusing on high-energy nuclear recoils would enhance the relative contribution from extragalactic particles.
Figure 5: Angular dependence of the dark-matter velocity distribution. Probability-density functions are shown for all particles (blue), the LSR direction (orange), and the anti-LSR direction (green). The vertical dashed red line indicates the average escape velocity of $520\ \mathrm{km\,s^{-1}}$.
@@
V DIRECT DETECTION EXPERIMENTS
@@
Figure 6: Extragalactic enhancement factor $\bigl{[}1+\Delta R/R\bigr{]}$ for the spin–independent differential rate, plotted as a function of recoil energy $E_{R}$. The four coloured curves correspond to WIMP masses of $5\ \mathrm{GeV}$ (blue), $15\ \mathrm{GeV}$ (orange), $50\ \mathrm{GeV}$ (green) and $200\ \mathrm{GeV}$ (red). Each curve rises sharply once $E_{R}$ exceeds the Milky-Way escape-velocity kinematic threshold for that mass, indicating how the unbound (extragalactic) component boosts the event rate at high recoil energies while leaving the low-energy spectrum essentially unchanged.
[…]
@@ VI Discussion and Implications
The Local Group environment contributes 26% by number and 38% by flux of the local dark matter.We find a strong directional anisotropy, with extragalactic fractions ranging from near zero to 71% across the sky. In addition, recent N-body simulations [6] suggest that even higher velocity particles ($>700$ km/s) originate from galactic substructure. This creates a multi-component velocity structure with moderate-velocity extragalactic particles (400–600 km/s) dominating the total flux and rare high-velocity articles from Milky-Way satellite like the Large Magellanic Cloud. Next-generation directionally-sensitive detectors could potentially isolate these different components through their distinct kinematic signatures.

### v2

@@ Abstract
We calculate the contribution of extragalactic dark matter to the local dark matter density and flux in the Milky Way. By analyzing the Galactic escape velocity as a function of direction, we establish a criterion for separating particles bound to the Milky Way from those originating in the Local Group environment. Our analysis finds that about 26% of local particles are unbound to the Galaxy and contribute about 38% of the total mass flux. The sky pattern is anisotropic. We provide rate and modulation predictions relevant for current and future direct-detection experiments.
@@ I Introduction
Dark matter constitutes about 85% of the matter content in our universe, yet its nature remains elusive. Past predictions for the detection rate of laboratory experiments focused on particles that are gravitationally-bound to the Milky Way [1, 2, 3, 4].Here, we show that a significant component originates from beyond our Galaxy. Inclusion of the extragalactic contribution could impact the quantitative interpretation of direct detection experiments. [5]
To identify extragalactic dark matter particles, we calculate the escape velocity as a function of direction across the sky. Particles exceeding this threshold cannot be gravitationally bound to the Milky Way and must have an extragalactic origin. This condition provides a method for distinguishing galactic from extragalactic particles; those moving faster than the local escape velocity in a given direction must originate beyond the Milky Way’s gravitational influence [6]
The extragalactic component primarily originates from the Local Group environment surrounding our Galaxy. The Local Group, dominated by the Milky Way, Andromeda and the intergalactic matter around them, contains of order $\sim 5\times 10^{12}M_{\odot}$ of mass, with the baryonic component representing approximately 17% of the total [7]. This intergalactic medium supplies high-velocity dark matter particles, potentially detectable in Earth-based experiments. Recent N-body simulations [8] have demonstrated that the highest velocity dark matter particles (v >> 600 km/s) in Milky Way-like halos predominantly originate from fast-moving substructure, particularly Large Magellanic Cloud analogs, rather than from the broader extragalactic environment. Hydrodynamical Local-Group simulations independently confirm this picture and find no evidence for a separate extragalactic high-speed component, with the local velocity distribution well described by a truncated Maxwellian tail [9].This finding suggests a multi-component environment of high-velocity dark matter sources that must be considered in detection experiments.
@@ II Galactic Escape Velocity
Figure 1: Minimum escape velocity as a function of direction in a spherical projection of Galactic coordinates within our model. The map displays significant angular variation, with the highest escape velocities (>700>700 km/s, yellow region) concentrated around galactic coordinates $(90^{\circ},0^{\circ})$ in the eastern hemisphere, while the lowest escape velocities (<400<400 km/s, dark purple) appear in the opposite direction. This anisotropy reflects the LSR motion and the asymmetric gravitational potential of the Milky Way.
@@ IV.1 Extragalactic Flux Analysis
Figure 4: The extragalactic dark matter flux in GeV/cm2/s ($\times 10^{6}$), shows significant angular variation with the highest flux values ($\sim 2\times 10^{7}$ GeV/cm2/s, yellow regions) concentrated in the western galactic hemisphere and a pronounced minimum in the opposite direcction.
@@ IV.2 Implications for Detection
Our analysis indicates that the extragalactic component contributes $38\%$ of the total mass flux while representing only $25\%$ of the particle number density. This kinematic distinction suggests that energy-selective detection methods could preferentially sample the extragalactic population, even without explicit directional sensitivity. Specifically, focusing on high-energy nuclear recoils would enhance the relative contribution from extragalactic particles.  [18]
[…]
@@
V Escape-speed uncertainty and extragalactic fractions
We varied the escape speed between 420 and 650 km s-1 and recomputed the sky-averaged fraction of particles classified as extragalactic using the criterion $v\geq v_{\rm esc}(\theta,\phi)$. The mean fractions are 51.1%, 43.4%, 36.1%, 27.4%, 21.7%, 16.9%, 11.7%, and 8.7% at $v_{\rm esc}=420,450,480,520,550,580,620,650$ km s-1, respectively. The maximum directional fractions across the sky over the same set are 70.3%, 61.4%, 52.4%, 40.9%, 33.1%, 26.2%, 18.5%, and 13.9%. The extragalactic fraction is therefore strongly anti-correlated with the assumed escape speed (Pearson $r\simeq-0.99$ for the mean values).
For the nominal $v_{\rm esc}=520$ km s-1 case, the sky-averaged extragalactic fraction is 27.4% and the maximum directional fraction is 40.9%. The total mass flux is largely unchanged by this sweep, while the extragalactic mass-flux component varies at the 30–40% level across the same range.
Figure 6: Effect of varying the escape speed on extragalactic dark matter fractions and mass flux. The panels show how key observables respond to uncertainties in the Galactic escape velocity, demonstrating the sensitivity of EGDM component identification to this fundamental parameter. Variations in $v_{\rm esc}$ directly affect the boundary between gravitationally bound and unbound populations, with systematic shifts observable in both the integrated fractions and directional flux distributions.
Figure 7: Extragalactic dark matter fraction maps across the tested escape-speed values. All-sky projections in Mollweide format illustrate how the spatial distribution of EGDM fractions varies with different assumptions for $v_{\rm esc}$. Higher escape velocities incorporate more high-velocity particles into the bound population, systematically reducing the EGDM fraction across all sky directions while preserving the overall angular structure. The consistent dipole-like pattern across all $v_{\rm esc}$ values indicates that the directional anisotropy is robust to escape velocity uncertainties.
VI Component separation and flux budget
We separate three kinematic components using the escape-speed field: bound Galactic ($v<v_{\rm esc}$), intermediate-speed extragalactic ($v_{\rm esc}\leq v<700$ km s-1), and high-speed substructure ($v\geq 700$ km s-1). For the nominal model the sky-averaged flux fractions are 66.4% (Galactic), 26.3% (extragalactic), and 7.3% (substructure). The corresponding sky-averaged mass fluxes are 4.839, 3.378, and $0.966\times 10^{6}$ GeV cm-2 s-1, yielding a total of $9.182\times 10^{6}$ GeV cm-2 s-1. The extragalactic-to-total mass-flux ratio is about 0.47 in this configuration.
The directional maximum of the extragalactic fraction is about 70%. The combined extragalactic-plus-substructure fraction averages 33.6% across the sky.
Figure 8: K-values Decomposition of dark matter velocity components and their corresponding mass flux distributions. Top row: Velocity component fractions showing the galactic component ($v<v_{\rm esc}$), diffuse EGDM component (vesc << v << 700 km/s), and bound substructure component (v >> 700 km/s). Bottom row: Mass flux distributions in units of 10-3 GeV cm-2 s-1 for EGDM and substructure components, along with their flux ratio. The maps reveal the complex angular dependence of dark matter flux arising from the superposition of multiple velocity components in the Galactic halo, with the EGDM/Substructure flux ratio showing spatial variations by factors of 2–6 across the sky.
Decomposition of dark matter velocity components and their corresponding mass flux distributions. Top row: Velocity component fractions showing the galactic component ($v<v_{\rm esc}$), diffuse EGDM component (vesc << v << 700 km/s), and bound substructure component (v >> 700 km/s). Bottom row: Mass flux distributions in units of 10-3 GeV cm-2 s-1 for EGDM and substructure components, along with their flux ratio. The maps reveal the complex angular dependence of dark matter flux arising from the superposition of multiple velocity components in the Galactic halo, with the EGDM/Substructure flux ratio showing spatial variations by factors of 2–6 across the sky.
VII Spherical-harmonic content of the flux
We expand the flux map in spherical harmonics up to $\ell=4$. Using our sampling, the monopole power is $P_{0}\simeq 9.5\times 10^{14}$ (arbitrary units), the dipole power is $P_{1}\simeq 1.05\times 10^{15}$, and the quadrupole power is $P_{2}\simeq 2.16\times 10^{14}$. The ratios $P_{1}/P_{0}\simeq 1.11$ and $P_{2}/P_{0}\simeq 0.23$ summarize the large-scale anisotropy in this model.
Figure 9: Dependence of mean EGDM fraction on Galactic escape velocity. Error bars represent the directional spread (standard deviation) of EGDM fractions across the sky, while the shaded band indicates an additional ±20.8% systematic uncertainty. The strong inverse correlation demonstrates the sensitivity of extragalactic component identification to $v_{\rm esc}$ measurements
VIII Detector sensitivity and sky coverage
We compare the predicted modulation against representative systematic floors: DAMA (0.030 cpd kg-1 keV-1), ANAIS and COSINE (0.020), a conservative floor (0.020), and an aggressive benchmark (0.010). With our count-rate scaling, the detectable sky fractions are about 30% (DAMA), 35% (ANAIS/COSINE/conservative), and 37% (aggressive). The maximum predicted modulation among detectable directions is about 0.13 cpd kg-1 keV-1. A year-scale toy simulation of daily counts yields a modulation significance of about $1.5\,\sigma$ for the nominal configuration.
Figure 10: Directional detectability maps for various dark matter direct detection experiments. Sky maps in Mollweide projection showing the detectability (1 = Yes, 0 = No) for five different experimental configurations. Blue regions indicate directions where dark matter signals are detectable, while red regions indicate directions below the detection threshold. The detectability patterns show a clear hemispherical asymmetry, with approximately 50–70% sky coverage depending on the experimental sensitivity. All experiments exhibit similar overall patterns, with detectability concentrated in one hemisphere, though subtle differences exist in the boundary regions between detectable and undetectable areas, reflecting varying detector sensitivities and background rejection capabilities.
IX Recoil-energy rate enhancement
We evaluate the differential-rate enhancement factor $1+\Delta R/R]$ by comparing a truncated Maxwellian baseline to a baseline plus an added high-speed component above the escape speed. For WIMP masses of 5, 15, 50, and 200 GeV, the maximal enhancements in the accessible recoil-energy range are approximately 1.88, 2.66, 1.49, and 1.00, respectively. The enhancement turns on once $v_{\min}(E_{R})$ approaches the escape-speed threshold for a given mass and remains close to unity at low recoil energies.
X DIRECT DETECTION EXPERIMENTS
@@
Figure 11: Extragalactic enhancement factor $\bigl[1+\Delta R/R\bigr]$ for the spin–independent differential rate, plotted as a function of recoil energy $E_{R}$. The four coloured curves correspond to WIMP masses of 5GeV5\ $\mathrm${GeV} (blue), 15GeV15\ $\mathrm${GeV} (orange), 50GeV50\ $\mathrm${GeV} (green) and 200GeV200\ $\mathrm${GeV} (red). Each curve rises sharply once $E_{R}$ exceeds the Milky-Way escape-velocity kinematic threshold for that mass, indicating how the unbound (extragalactic) component boosts the event rate at high recoil energies while leaving the low-energy spectrum essentially unchanged.
Because the Earth’s orbital velocity ($v_{\oplus}\simeq 30\ \text{km\,s}^{-1}$) is comparable to the uncertainty in the escape-velocity threshold (Sec. II), the extragalactic flux exhibits an annual modulation with amplitude $A\equiv(R_{\max}-R_{\min})/(R_{\max}+R_{\min})\approx 0.33$ for the fiducial Germanium detector. This is an order-of-magnitude larger than the 6–7% modulation predicted in the Standard Halo Model (SHM) and is therefore a distinctive signature testable by modulation searches such as ANAIS [21] and COSINE-100 [22]. For a 250kg250\ $\text${kg} NaI(Tl) array operated for 7 yr, our flux enhancement would raise the expected modulation significance from $1.8\sigma$ (SHM only) to $3\sigma$.
@@ XI Discussion and Implications
The Local Group environment contributes 26% by number and 38% by flux of the local dark matter.We find a strong directional anisotropy, with extragalactic fractions ranging from near zero to 71% across the sky. In addition, recent N-body simulations [8] suggest that even higher velocity particles (>700>700 km/s) originate from galactic substructure. This creates a multi-component velocity structure with moderate-velocity extragalactic particles (400–600 km/s) dominating the total flux and rare high-velocity articles from Milky-Way satellite like the Large Magellanic Cloud. Next-generation directionally-sensitive detectors could potentially isolate these different components through their distinct kinematic signatures.